Tomorrow morning around 8:30, the hall will be opened and surveyed in preparation for the Restricted Access to install the RICH. Around 9am the upload of the Hall C viewer-software will take place.
During swing and owl shift, the RICH will be flushed, and we'll continue to take 'production' data on the waterfall target to exercise the DAQ, optimize cuts for PID, and understand the trigger efficiency.
Hopefully, by Friday day the RICH will be completely ready and we'll begin taking production data with the RICH on the waterfall target.
Special Instructions: --------------------- - The fast-feedback has been turning itself off, so keep on eye on its status. Both position and energy (RF) feedback should be ON. - Watch the energy-spread calculation from the SLI. The energy-spread should be smaller than 3e-5 (sigma), though they have been having problems with this at the higher currents.
If in doubt, check the experiment web-page or call the Run Coordinator.
A copy of this log entry has been emailed to: e94-107@jlab.org